OpenStudy (anonymous):

What is the radius of a sphere if its surface area is 2,122.64 square inches? (Use 3.14 for π.)

OpenStudy (anonymous):

We want to solve for r, so we can do the following: \[4\pi r^2\]=2122.64 Divide by 4: \[\pi r^2\]=530.66 Divide by 3.14: \[r^2\]=169 \[\sqrt{r^2}=\sqrt{169}\] r=13